OpenStudy (littleheron):

which of the following is not a perfect square? A. 4 B. 300 C. 81 D. 36

OpenStudy (witty_roger):

300

OpenStudy (solomonzelman):

\(\color{#0000ff }{ \displaystyle 16=4\cdot 4=4^2 }\) So, \(16\) is a perfect square, because you are multiplying a number \(4\) times itself (also \(4\)) to obtain \(16\)\(\bf .\) \(\color{#0000ff }{ \displaystyle 25=5\cdot 5=5^2 }\) So, \(25\) is a perfect square, because you are multiplying a number \(5\) times itself (also \(5\)) to obtain \(25\)\(\bf .\) \(\color{#0000ff }{ \displaystyle 1=1\cdot 1=1^2 }\) So, \(1\) is a perfect square, because you are multiplying a number \(1\) times itself (also \(1\)) to obtain \(1\)\(\bf .\) \(\color{#ff0000 }{ \displaystyle 0=0\cdot 0=0^2 }\) That is true, BUT... THIS IS AN EXCEPTION, \(0\) IS NOT A PERFECT SQUARE! \(\color{#0000ff }{ \displaystyle 9=3\cdot 3=3^2 }\) So, \(9\) is a perfect square, because you are multiplying a number \(3\) times itself (also \(3\)) to obtain \(9\)\(\bf .\)

OpenStudy (littleheron):

soo... 81?

OpenStudy (solomonzelman):

81 = 9 × 9 = 9\(^2\)

OpenStudy (solomonzelman):

So \(81\) is a perfect square

OpenStudy (littleheron):

then 300

OpenStudy (solomonzelman):

yes, 300 is NOT a perfect square.

OpenStudy (solomonzelman):

You can make the following definition \(\small(\)below\(\small) \). Suppose there is a natural number \(\color{#0000ff }{ \displaystyle \ell }\). So, if \(\color{#0000ff }{ \displaystyle \ell \times \ell=x}\) Then, \(\color{#0000ff }{ \displaystyle x}\) is a perfect square.

OpenStudy (witty_roger):

medal plz @LittleHeron #Being_selfish

OpenStudy (solomonzelman):

The main thing is the learning, not medals... According to the policy, you should not: - Give direct answers - Ask for medals explicitely.

OpenStudy (witty_roger):

hmmm

OpenStudy (witty_roger):

@littleheron wud give me one if he wants to